diameter of sphere is decreased by 25% by what percent does its curved surface area decreases
Answer by Alan3354(69443)

SA is a function of the square of the diameter.
d2 = 0.75*d1
SA2 = 0.75^2 * SA1 = 0.5625SA1
A decrease of 43.75%